Solutions of sodium carbonate and silver nitrate react to form solid silver carbonate and a solution of sodium nitrate. A solution containing 3.25 g of sodium carbonate is mixed with one containing 6.50 g of silver nitrate. After the reaction is complete, the solutions are evaporated to dryness, leaving a mixture of salts. How many grams of each of the following compounds are present after the reaction is complete?


sodium carbonate


silver nitrate


silver carbonate


sodium nitrate


1
Expert's answer
2021-02-10T01:46:31-0500

Na2CO3 + 2 AgNO3 "\to" AgCO3 "\downarrow" + 2NaNO3

n(Na2CO3) = 3.25g/106g/mol = 0.0307mol

n(AgNO3) = 6.50g/170g/mol = 0.0382mol

1mol Na2CO3 "\to" 2mol AgNO3

xmol Na2CO3 "\to" 0.0382 mol AgNO3

x = 0.0191

n(Na2CO3) = 0.0307 - 0.0191 = 0.0116mol

m(Na2CO3) = 0.0116mol × 106 g/mol = 1.2296g

n(NaNO3) = 0.0382mol

m(NaNO3) = 0.0382mol × 85g/mol = 3.247 g

n(Ag2CO3) = 0.0191 mol

m(Ag2CO3) = 0.0191mol × 276 g/mol = 5.2716g

Sodium carbonate = 1.2296g

Silver nitrate = 0g

Silver carbonate = 5.2716g

Sodium nitrate = 3.247g